# Modified gravitational wave propagations in linearized gravity with Lorentz and diffeomorphism violations and their gravitational wave constraints

### v1

@@ IV Bayesian inference and parameter estimation
where denotes the complex conjugate, and $S(f)$ is the power spectral density (PSD) of the detector. To ensure stable and reliable parameter estimation, we use the PSD data provided in the LVK posterior sample, which is more robust compared to PSDs derived from strain data using Welch averaging Abbott et al. (2019a); Cornish and Littenberg (2015); Littenberg and Cornish (2015).
@@
Considering that the spherical expansion coefficient formulae in Eqs. (2.49, 2.47, 2.48) provide a general solution for different events within the same coordinate system, we can combine the posterior distributions of individual events as:
@@ V Results
In this section, we present the results of the constraints on the Lorentz- and diffeomorphism-violating coefficients for mass diemsnion $d=2$ and $d=3$ cases. In the following, we present the results for $d=2$ and $d=3$ separately.
@@ V.1 $d=2$
Here we adopt a Planck cosmology with $\Omega_{m}=0.315$, $\Omega_{\Lambda}=0.685$, and $H_{0}=67.4\;{\rm km}\;{\rm s}^{-1}\;{\rm Mpc}^{-1}$ Aghanim et al. (2020). The parameter $A_{\bar{\mu}}$ is the parameter we sampled in the Bayesian inference along with other GR parameters. We use the uniform prior for parameter $A_{\bar{\mu}}$ in our analysis. Then from the marginal posterior distributions of $A_{\bar{\mu}}$ and the redshift $z$ of the analyzed GW events, one can obtain posterior distributions of $k_{(I)00}^{(2)}$. In Fig. 1 we display the marginalized posterior distributions of $k_{(I)00}^{(2)}$ from selected GW events in the GWTC-3. For most GW events we analyze in each test, we do not find any significant signatures of Lorentz and diffeomorphism violation due to the coefficient $k_{(I)00}^{(2)}$. A few events that suggest nonzero values for the non-GR coefficients $A_{\bar{\mu}}$ are excluded from our analysis due to their contradiction with GR. The posterior posterior distributions for $k_{(I)00}^{(2)}$ from the excluded GW events are presented in Fig. 2. It is mentioned in Wang (2020); Wang et al. (2022) that these results may arise from limitations in current waveform approximants, such as systematic errors during the merger phase or unaccounted physical effects like eccentricity. Consequently, we have excluded these events from our analysis. Table 2 presents the list of excluded events along with the estimated constraints on $k^{(2)}_{(I)00}$ for each excluded event.
@@ V.2 $d=3$
Then the posterior samples of each component of $k_{(V)jm}^{(3)}$ can be calculated from the marginal posterior distributions of $A_{\mu}$, right ascension (ra), declination (dec), and redshift $z$ of the analyzed GW events, via Eqs. (5.5). Fig. 3 presents the marginalized posterior distributions of $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$ from selected GW events in the GWTC-3. The lower and upper limits of each coefficient $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$. $k_{(I)jm}^{(6)}$ are represented by the vertical dash line in each figure of Fig. 3. For most GW events we analyze in each test, we do not find any significant signatures of Lorentz and diffeomorphism violation due to the coefficient $k_{(V)ij}^{(3)}$. In addition, for each coefficient of $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$, we consider each of them as a universal parameter for all GW events, and then we obtain their combined constraints separately by multiplying the posterior distributions of the individual events together. Table. 3 summarizes the $90\%$ confidence interval of each coefficients $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$. From both the Fig. 3 and Table. 3, it is obvious that the posterior samples and the $90\%$ confidence interval of each coefficient $k_{(V)ij}^{(3)}$ are all consistent with zero, which indicates there are no any signatures of the Lorentz and diffeomorphism violations arising in the linearized gravity of SME has been found in the GW signals.
@@
Figure 3: The posterior distributions for $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$ from selected GW events in the LIGO-Virgo-KAGRA catalog GWTC-3. The legend indicates the events that give the six tightest constraints. The vertical dash line denotes the 90% interval for the coefficients $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$ from combined results.
@@ Acknowledgements
We appreciate the helpful discussions with Bo-Yang Zhang and Yuan-Zhu Wang. This work is supported by the National Key Research and Development Program of China under Grant No. 2020YFC2201503, the National Natural Science Foundation of China under Grants No. 12275238 and No. 11675143, the Zhejiang Provincial Natural Science Foundation of China under Grants No. LR21A050001 and No. LY20A050002, and the Fundamental Research Funds for the Provincial Universities of Zhejiang in China under Grant No. RF-A2019015. W.Z. is supported by the National Natural Science Foundation of China (Grants No. 12325301 and 12273035) and the National Key Research and Development Program of China (Grant No. 2022YFC2204602 and 2021YFC2203102), the Fundamental Research Funds for the Central Universities.

### v2

@@ IV Bayesian inference and parameter estimation
where denotes the complex conjugate, and $A(f)$ represents the measured GW strain signal at the LIGO/Virgo/KAGRA detectors, and $B(f)$ represents the theoretical GW predicted by the waveform model. $S(f)$ is the power spectral density (PSD) of the detector. To ensure stable and reliable parameter estimation, we use the PSD data provided in the LVK posterior sample, which is more robust compared to PSDs derived from strain data using Welch averaging Abbott et al. (2019a); Cornish and Littenberg (2015); Littenberg and Cornish (2015).
@@
Considering that the spherical expansion coefficient formulae in Eqs. (2.76, 2.74, 2.75) provide a general solution for different events within the same coordinate system, we can combine the posterior distributions of individual events as:
@@ V Results
In this section, we present the results of the constraints on the Lorentz- and diffeomorphism-violating coefficients for mass dimensions $d=2$ and $d=3$ cases. In the following, we present the results for $d=2$ and $d=3$ separately.
@@ V.1 $d=2$
Here we adopt a Planck cosmology with $\Omega_{m}=0.315$, $\Omega_{\Lambda}=0.685$, and $H_{0}={1.44\times 10^{-42}\;{\rm GeV}}$ Aghanim et al. (2020). The parameter $A_{\bar{\mu}}$ is the parameter we sampled in the Bayesian inference along with other GR parameters. We use the uniform prior for parameter $A_{\bar{\mu}}$ in our analysis. Then from the marginal posterior distributions of $A_{\bar{\mu}}$ and the redshift $z$ of the analyzed GW events, one can obtain posterior distributions of $k_{(I)00}^{(2)}$. In Fig. 1 we display the marginalized posterior distributions of $k_{(I)00}^{(2)}$ from selected GW events in the GWTC-3. For most GW events we analyze in each test, we do not find any significant signatures of Lorentz and diffeomorphism violation due to the coefficient $k_{(I)00}^{(2)}$. A few events that suggest nonzero values for the non-GR coefficients $A_{\bar{\mu}}$ are excluded from our analysis due to their contradiction with GR. The posterior posterior distributions for $k_{(I)00}^{(2)}$ from the excluded GW events are presented in Fig. 2. It is mentioned in Wang (2020); Wang et al. (2022) that these results may arise from limitations in current waveform approximants, such as systematic errors during the merger phase or unaccounted physical effects like eccentricity. Consequently, we have excluded these events from our analysis. Table 2 presents the list of excluded events along with the estimated constraints on $k^{(2)}_{(I)00}$ for each excluded event.
@@ V.2 $d=3$
Then the posterior samples of each component of $k_{(V)jm}^{(3)}$ can be calculated from the marginal posterior distributions of $A_{\mu}$, right ascension (ra), declination (dec), and redshift $z$ of the analyzed GW events, via Eqs. (5.5). Fig. 3 presents the marginalized posterior distributions of $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$ from selected GW events in the GWTC-3. The lower and upper limits of each coefficient $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$. $k_{(I)jm}^{(6)}$ are represented by the vertical dash line in each figure of Fig. 3. For most GW events we analyze in each test, we do not find any significant signatures of Lorentz and diffeomorphism violation due to the coefficient $k_{(V)ij}^{(3)}$. In addition, for each coefficient of $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$, we consider each of them as a universal parameter for all GW events, and then we obtain their combined constraints separately by multiplying the posterior distributions of the individual events together. Table. 3 summarizes the $90\%$ confidence interval of each coefficients $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$. From both the Fig. 3 and Table. 3, it is obvious that the posterior samples and the $90\%$ confidence interval of each coefficient $k_{(V)ij}^{(3)}$ are all consistent with zero, which indicates there are no signatures of the Lorentz and diffeomorphism violations arising in the linearized gravity of SME has been found in the GW signals.
@@
Figure 3: The posterior distributions for $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$, and ${\rm Im}k_{(V)11}^{(3)}$ from selected GW events in the LIGO-Virgo-KAGRA catalog GWTC-3. The legend indicates the events that give the six tightest constraints. The vertical dash lines denote the 90% intervals for the coefficients $k_{(V)00}^{(3)}$, $k_{(V)10}^{(3)}$, ${\rm Re}k_{(V)11}^{(3)}$ from combined results.
@@ Acknowledgements
We appreciate the helpful discussions with Dr. Xiao Zhi, Bo-Yang Zhang, and Yuan-Zhu Wang. This work is supported by the National Key Research and Development Program of China under Grant No. 2020YFC2201503, the National Natural Science Foundation of China under Grants No. 12275238 and No. 11675143, the Zhejiang Provincial Natural Science Foundation of China under Grants No. LR21A050001 and No. LY20A050002, and the Fundamental Research Funds for the Provincial Universities of Zhejiang in China under Grant No. RF-A2019015. W.Z. is supported by the National Natural Science Foundation of China (Grants No. 12325301 and 12273035) and the National Key Research and Development Program of China (Grant No. 2022YFC2204602 and 2021YFC2203102), the Fundamental Research Funds for the Central Universities.